Hi! I could use some help with a word problem I have!
The question is:
Judy, who is 17 years old, has a grandmother 77 years. How many years ago was Judy's grandmother 6 times as old as Judy?
I tried writing things like:
j=17
G=77
I also multiplied 17 by 6, but I got 102, which would make grandmother even older than what she is now, so I thought that made no sense.
Could you explain this to me? Thank you so much!

Answer by lwsshak3(11628)

Judy, who is 17 years old, has a grandmother 77 years. How many years ago was Judy's grandmother 6 times as old as Judy?
let x=years ago when grandmother was 6 times as old as judy
17-x=judy's age x yrs ago
77-x=grandmother's age x yrs ago
..
77-x=6(17-x)
77-x=102-6x
5x=25
x=5
years ago when grandmother was 6 times as old as judy=5